A thin circular ring of area A is held perpendicular to a uniform magnetic field of induction B. A small cut is made in the ring and a galvanometer is connected across the ends such that the total resistance of the circuit is R. When the ring is suddenly squeezed to zero area, the charge flowing through the galvanometer is.

Grade:11

1 Answers

Kevin Nash
9 years ago
Hello Student,
Please find the answer to your question
(b) The current induced will be
i = |e| / R ⇒ i = 1/ R dϕ/dt But i = dq / dt
⇒ dq / dt = 1/R dϕ / dt ⇒ ∫ dq = 1/R ∫ dϕ⇒ q = BA/R
Thanks
Kevin Nash
askIITians Faculty